OpenStudy (anonymous):

Determine if triangle DEF with coordinates D (3, 2), E (4, 6), and F (7, 3) is an equilateral triangle. Use evidence to support your claim. If it is not an equilateral triangle, what changes can be made to make it equilateral? Be specific.

OpenStudy (anonymous):

I used the distance formula for each segment in triangle DEF and have come to the conclusion that the triangle is not an equilateral triangle. The distance for Segments DE and DF are sq.rt.17, but the distance for Segment EF is sq.rt. 18. I am having difficulty answering the second part of the question, "If it is not an equilateral triangle, what changes can be made to make it equilateral? Be specific." Please provide guidance as to how to go about answering this question. Thank you!

OpenStudy (anonymous):

Find the distance between D and E to determine the length of Segment DE. D = (3,2) and E = (4,6) Step 1: d = sq.rt. (x2 - x1)^2 + (y2 - y1)^2 Step 2: d = sq.rt. (4 - 3)^2 + (6 - 2)^2 Step 3: d = sq.rt. (1)^2 + (4)^2 Step 4: d = sq.rt. 1 + 16 Step 5: d = sq.rt. 17 Find the distance between E and F to determine the length of Segment EF. E = (4,6) and F = (7,3) Step 1: d = sq.rt. (x2 - x1)^2 + (y2 - y1)^2 Step 2: d = sq.rt. (7 - 4)^2 + (3 - 6)^2 Step 3: d = sq.rt. (3)^2 + (-3)^2 Step 4: d = sq.rt. 9 + 9 Step 5: d = sq.rt. 18 Find the distance between D and F to determine the length of Segment DF. Step 1: d = sq.rt. (x2 - x1)^2 + (y2 - y1)^2 Step 2: d = sq.rt. (7 - 3)^2 + (3 - 2)^2 Step 3: d = sq.rt. (4)^2 + (1)^2 Step 4: d = sq.rt 16 + 1 Step 5: d = sq.rt. 17

OpenStudy (danjs):

i guess, sqrt(18) - sqrt(17) needs to be taken off EF, and recalculate the position of one of the points

OpenStudy (anonymous):

I agree, but I have tried different combinations and have failed to find an alternative point.

OpenStudy (loser66):

By distance formula , you have \(EF =\sqrt{(7-4)^2+(3-6)^2}=\sqrt {18}\) Only one way you can do is to make it down to \(\sqrt{17}\)

OpenStudy (anonymous):

That is what I thought before, but I got points off because I didn't give an alternative point.

OpenStudy (loser66):

by changing the location of E. F cannot be changed because it relates to D

OpenStudy (loser66):

if E( 6, 7) then you get the required triangle.

OpenStudy (anonymous):

Seriously!?!

OpenStudy (loser66):

Read the problem, they say : by what changes??

OpenStudy (anonymous):

Yes, thank you so much!

OpenStudy (loser66):

my change is the point E from (4, 6) to (6,7)

OpenStudy (danjs):

E or F can move in a circle around point D, right

OpenStudy (loser66):

@DanJS It is a good idea but I don't think they use the orbit concept here.

OpenStudy (anonymous):

I plugged in 6,7 yet still got different measures for each distance. Did I do something wrong?

OpenStudy (loser66):

oh, yea!! I am wrong!!

OpenStudy (loser66):

since if you change E, then DE changes also. OMG. I am terribly sorry.

OpenStudy (anonymous):

No worries!!

OpenStudy (danjs):

I dont see how that has to happen, if you move E , DE can stay the same length, the radius of the circle, when you move E on an Arc towards F

OpenStudy (danjs):

or F towards E the same way

OpenStudy (anonymous):

Circle?

OpenStudy (loser66):

I know how to do!! Thanks God. He gives me a chance to fix my mistake. hehehe.

OpenStudy (anonymous):

Lol thank you both for your help and patience!

OpenStudy (danjs):

Circle centered at D, radius root(17), you just swing F towards E or E towards F, till the chord EF is root17

OpenStudy (danjs):

i am guessing

OpenStudy (danjs):

or set up two distance formula , and use (x,y) for F, and figure out x and y, possibly

OpenStudy (loser66):

OK. \(DF=\sqrt{17}\)

OpenStudy (anonymous):

I tried your second option, but every time I changed a coordinate, I failed to get the same measure for each segment.

OpenStudy (loser66):

\(DE=\sqrt{(x_E-3)^2+(y_E-2)^2)}\) \(EF=\sqrt{(7-x_E)^2+(3-y_E)^2)}\)

OpenStudy (loser66):

and we want DE = EF, so, solve for it!!

OpenStudy (loser66):

\(DE = EF = \sqrt{17}\)

OpenStudy (danjs):

that what i said.. :)

OpenStudy (loser66):

ok, take over, please. hehehe.. I tried to fix my mistake but if you know how to do, please, finish the stuff. I am not good at explaining. Please :)

OpenStudy (anonymous):

So the change or different coordinate is (x,y)?

OpenStudy (danjs):

there are two possible points for F, one on each side of the Y axis, i think

OpenStudy (danjs):

17 = (x-4)^2 + (y-6)^2 17 = (x-3)^2 + (y-2)^2 i solved those with a computer,

OpenStudy (anonymous):

So D still equals (3,2)...E still equals (4,6). What does F = ?

OpenStudy (danjs):

\[F(x,y) = (\frac{ 4\sqrt{3} + 7 }{ 2 } , \frac{ 8-\sqrt{3} }{ 2 }) \approx (6.9641, 3.1340)\]

OpenStudy (anonymous):

Woah...that is very precise. I think I may have gone about answering this question the wrong way because I doubt my teacher would have expected such an answer.

OpenStudy (anonymous):

Do you think I should answer using your original suggestion, "i guess, sqrt(18) - sqrt(17) needs to be taken off EF, and recalculate the position of one of the points,"....?

OpenStudy (danjs):

\[FD = FE = \sqrt{17}\] that what those 2 distance equations above are

OpenStudy (anonymous):

Thank you for your help and time! If you could, would you please help me formulate my final answer. If not, I totally understand because I have already taken up so much of your time!!

OpenStudy (danjs):

draw a picture first

OpenStudy (danjs):

label everything

OpenStudy (danjs):

then maybe, move F up a little, and relable that F' point (x,y), EF' = root (17) to make an equilateral triangle

OpenStudy (danjs):

then use the work above

OpenStudy (anonymous):

Okay, that is perfect! Thanks again. Can I give you multiple awards...other than just a medal?

OpenStudy (anonymous):

*other not multiple

OpenStudy (danjs):

only things i know are fan and testimonial

OpenStudy (anonymous):

Testimonial?

OpenStudy (danjs):

@Loser66 did just as much...

OpenStudy (danjs):

hover over name, and buttons are there

OpenStudy (anonymous):

Okay, I gave you a medal, became a fan and submitted a testimonial. Thanks!!

OpenStudy (danjs):

hah, thanks, ill fan you to

OpenStudy (anonymous):

No need!! Good night!!